solve by elimination method
9x-9y=9
7y-3x=-5

Answer by checkley77(12844)

9x-9y=9
7y-3x=-5 MULTIPLY BY 3 & ADD
-9X+21Y=-15
9X-9Y=9
------------------
3Y=-6
Y=-6/3
Y=-2 ANS.
9X-9*-2=9
9X+18=9
9X=9-18
9X=-9
X=-9/9
X=-1 ANS.